The average annual tuition and fees at all 4-year institutions in the US in 1982 was $10,385 and in 2012 was $ 23,872. Let y be the average tuition and fees in the year x, where x = 0 represents the year 1982.
a) Write a linear equation, in slope-intercept form, that models the growth in average tuition and fees at all 4-year institutions in the US in terms of the year x.











b) Use this equation to predict the average tuition and fees at 4-year institutions in the US in the year 2030.









c) Explain what the slope of this line means in the context of the problem.
Please show all work.

Answer by stanbon(75887)

The average annual tuition and fees at all 4-year institutions in the US in 1982 was $10,385 and in 2012 was $ 23,872. Let y be the average tuition and fees in the year x, where x = 0 represents the year 1982.
a) Write a linear equation, in slope-intercept form, that models the growth in average tuition and fees at all 4-year institutions in the US in terms of the year x.
------
Let 1982 be 0
Then 2012 is 30
---------------------
You have 2 points:: (0,10385) and (30,23872)
----
slope = (23872-10385)/30 = 449.57
intercept = 10385
-----------------------
Equation:
y = 449.57x + 10385
----

b) Use this equation to predict the average tuition and fees at 4-year institutions in the US in the year 2030.
2030 would be 2030-1982 = 48
---
f(48) = 449.47*48 + 10385 = 31964
----------------------
c) Explain what the slope of this line means in the context of the problem.
Please show all work.
Each year the tuition increases $449.57
-----------
Cheers,
Stan H.
-----------